# LLM-Powered GUI Agents in Phone Automation: Surveying Progress and Prospects

### v1

@@ 6 Challenges and Future Directions
Lightweight and Efficient On-Device Deployment. Deploying LLMs on mobile devices confronts substantial computational and memory constraints. Current hardware often struggles to support large models with minimal latency and power consumption. Approaches such as model pruning, quantization, and efficient transformer architectures can address these constraints [100]. Recent innovations demonstrate promising progress. Octopus v2 [265] shows that a 2-billion parameter on-device model can outpace GPT-4 in accuracy and latency, while Lightweight Neural App Control [75] achieves substantial speed and accuracy improvements by distributing tasks efficiently. Moreover, specialized hardware accelerators and edge computing solutions can further reduce dependency on the cloud, enhance privacy, and improve responsiveness [52]. Consider leveraging the powerful code generation capabilities of small language models (SLMs) to transform GUI task automation into a code generation problem. This approach fully utilizes the strengths of SLMs, significantly enhancing the efficiency and performance of GUI agents on mobile devices [266, 267].
User-Centric Adaptation: Interaction and Personalization. Current agents often rely on extensive human intervention to correct errors or guide task execution, undermining seamless user experiences. Enhancing the agent’s ability to understand user intent and reducing manual adjustments is crucial. Future research should improve natural language understanding, incorporate voice commands and gestures, and enable agents to learn continuously from user feedback [61, 51, 52, 268]. Personalization is equally important. One-size-fits-all solutions are insufficient given users’ diverse preferences and usage patterns. Agents should quickly adapt to new tasks and user-specific contexts without costly retraining. Integrating manual teaching, zero-shot learning, and few-shot learning can help agents generalize from minimal user input [73, 61, 79, 269], making them more flexible and universally applicable. For example, AdaptAgent [205] is capable of adapting to entirely new domains with as few as two human demonstrations. This not only proves the efficiency of limited human input, but also paves a new path for the development of multi-modal agents with broad adaptability.
@@
Ensuring Reliability and Security. As agents gain access to sensitive data and perform critical tasks, reliability and security are paramount. Current systems may be susceptible to adversarial attacks, data breaches, and unintended actions [270]. At the same time, LLM agents are also susceptible to backdoor attacks [271, 272]. Robust security protocols, error-handling techniques, and privacy-preserving methods are needed to protect user information and maintain user trust [67, 103]. Employing techniques such as data localization, encrypted communication, and anonymization can effectively protect user privacy while collecting data [262]. Continuous monitoring and validation processes can detect vulnerabilities and mitigate risks in real-time [61]. Ensuring that agents behave predictably, respect user privacy, and maintain consistent performance under challenging conditions will be crucial for widespread adoption and long-term sustainability.

### v2

@@ 5.2.1 Evaluation Pipelines
LearnGUI [155] offers a novel approach by introducing the first comprehensive benchmark specifically designed for demonstration-based learning in mobile GUI agents. Rather than pursuing universal generalization through larger datasets, it focuses on improving agent performance in unseen scenarios through human demonstrations. The benchmark comprises 2,252 offline tasks and 101 online tasks with high-quality human demonstrations.
@@
AutoEval [154] addresses the practicality and scalability challenges in mobile agent evaluation by introducing a framework that requires no manual effort to define task reward signals or implement evaluation codes. It employs a Structured Substate Representation to describe UI state changes during agent execution and utilizes a Judge System that can autonomously evaluate agent performance with over 94% accuracy compared to human verification.
@@ 6 Challenges and Future Directions
Lightweight and Efficient On-Device Deployment. Deploying LLMs on mobile devices confronts substantial computational and memory constraints. Current hardware often struggles to support large models with minimal latency and power consumption. Approaches such as model pruning, quantization, and efficient transformer architectures can address these constraints [111]. Recent innovations demonstrate promising progress. Octopus v2 [286] shows that a 2-billion parameter on-device model can outpace GPT-4 in accuracy and latency, while Lightweight Neural App Control [81] achieves substantial speed and accuracy improvements by distributing tasks efficiently. AppVLM [113], a lightweight vision-language model, matches GPT-4o in online task completion success rate while being up to ten times faster, making it practical for real-world deployment. Moreover, specialized hardware accelerators and edge computing solutions can further reduce dependency on the cloud, enhance privacy, and improve responsiveness [52]. Consider leveraging the powerful code generation capabilities of small language models (SLMs) to transform GUI task automation into a code generation problem. This approach fully utilizes the strengths of SLMs, significantly enhancing the efficiency and performance of GUI agents on mobile devices [287, 288].
User-Centric Adaptation: Interaction and Personalization. Current agents often rely on extensive human intervention to correct errors or guide task execution, undermining seamless user experiences. Enhancing the agent’s ability to understand user intent and reducing manual adjustments is crucial. Future research should improve natural language understanding, incorporate voice commands and gestures, and enable agents to learn continuously from user feedback [61, 51, 52, 289]. Personalization is equally important. One-size-fits-all solutions are insufficient given users’ diverse preferences and usage patterns. Agents should quickly adapt to new tasks and user-specific contexts without costly retraining. Integrating manual teaching, zero-shot learning, and few-shot learning can help agents generalize from minimal user input [79, 61, 85, 290], making them more flexible and universally applicable. For example, AdaptAgent [225] is capable of adapting to entirely new domains with as few as two human demonstrations. This not only proves the efficiency of limited human input, but also paves a new path for the development of multi-modal agents with broad adaptability. Similarly, LearnAct [155] demonstrates the power of human demonstrations in mobile GUI agents, using a multi-agent framework to automatically extract knowledge from demonstrations to enhance task completion. It establishes demonstration-based learning as a promising direction for creating more personalized and adaptive mobile agents.
@@
Ensuring Reliability and Security. As agents gain access to sensitive data and perform critical tasks, reliability and security are paramount. Current systems may be susceptible to adversarial attacks, data breaches, and unintended actions [291]. At the same time, LLM agents are also susceptible to backdoor attacks [292, 293]. Recent research like AEIA-MN [294] has demonstrated that multimodal LLM-powered mobile agents are highly vulnerable to Active Environmental Injection Attacks, where attackers manipulate environmental elements (e.g., notifications) to mislead agents, achieving attack success rates up to 93% in benchmark tests. Robust security protocols, error-handling techniques, and privacy-preserving methods are needed to protect user information and maintain user trust [71, 116]. Employing techniques such as data localization, encrypted communication, and anonymization can effectively protect user privacy while collecting data [283]. FedMABench [153] addresses the challenges of distributed training using federated learning, providing a comprehensive benchmark for evaluating mobile agents across heterogeneous environments. Continuous monitoring and validation processes can detect vulnerabilities and mitigate risks in real-time [61]. Ensuring that agents behave predictably, respect user privacy, and maintain consistent performance under challenging conditions will be crucial for widespread adoption and long-term sustainability.